OpenStudy (anonymous):

this is a simple question, 2x+y=4 how dose that turn into y=2x-4 in other words how did the 2x become positive? like 2x-4y=5 how did it tun into -4y=2x-5 how did 2x become positive?

OpenStudy (anonymous):

if u move the number or varialbe to other side the sign changes like example 2x-y=0 and u move y to other side than the y varialbe becomes positive which will be now 2x=y

OpenStudy (cggurumanjunath):

see the rule is that when numbers or terms are taken from one hand side (left or right) from the equal sign the term changes its sign.very important note.

OpenStudy (anonymous):

it usually depends on the quesiton if ur finding x or y variable than you need to make the variable alone and move numbers to others side or other variables

OpenStudy (anonymous):

and btw when its 2x+y=4 its gogin to be y=-2x+4 not positive

OpenStudy (anonymous):

the sign has to change when u move to other side

OpenStudy (anonymous):

any questions?

OpenStudy (anonymous):

no thanks for clearing that up!

OpenStudy (anonymous):

did i answer ur question?

OpenStudy (anonymous):

well the two equations above are suppose to be perpendicular but the slope is not negative reciprocals... so I thought i did the problem wrong @jkim3592

OpenStudy (anonymous):

oh thats the slope?

OpenStudy (anonymous):

when ur finding perpendicular slopes the all the signs change

OpenStudy (anonymous):

srry g2g

OpenStudy (anonymous):

y=-2x + 4, and y= 2/4x - 5/4 .the second is opposite sign and 2/4 = 1/2 is reciprocal to 2. i found equations applying each consecutive operation on both sides